40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Mar 6 12:34:30 2017, johannes, Summary, ASS, ASS light injection scenarios 
    Reply  Mon Aug 21 20:11:25 2017, johannes, Summary, General, Loss measurements plan 
       Reply  Mon Aug 21 21:26:41 2017, gautam, Summary, General, Loss measurements plan 
          Reply  Tue Aug 22 16:56:54 2017, johannes, Summary, General, AS laser existing components inventory 
             Reply  Mon Aug 28 23:22:56 2017, johannes, Update, PSL, PSL table auxiliary NPRO aom_switching_zoomin.pngaom_switching_zoomout.pngaom_switching_mini_zoomin.pngaom_switching_mini_zoomout.png
                Reply  Tue Aug 29 20:04:09 2017, rana, Update, PSL, PSL table auxiliary NPRO 
                   Reply  Tue Aug 29 21:36:59 2017, johannes, Update, PSL, PSL table auxiliary NPRO isomet_mini_1st_zoomout.pngisomet_mini_0th_zoomout.png
                      Reply  Tue Sep 5 23:13:44 2017, johannes, Update, PSL, PSL table auxiliary NPRO PSL_IR.pdfpsl_aux_laser.pdf
                         Reply  Thu Sep 7 23:09:00 2017, johannes, Update, PSL, PSL table auxiliary NPRO 3200-1113_200_switching_PD10A_zoomin.png3200-1113_200_switching_PD10A_zoomout.png
                            Reply  Mon Sep 11 12:40:32 2017, johannes, Update, PSL, PSL table auxiliary NPRO psl_aux_laser2.svg.pngPSL_IR2.png
Message ID: 13298     Entry time: Tue Sep 5 23:13:44 2017     In reply to: 13271     Reply to this: 13301
Author: johannes 
Type: Update 
Category: PSL 
Subject: PSL table auxiliary NPRO 

I used Gautam's mode measurement of the auxiliary NPRO (w=127.3um, z=82mm) for the spacing of the optics on the PSL table for the fiber injection and light modulation. As mentioned in previous posts, for the time being there is no Faraday isolator and no broadband EOM installed, but they're accounted for in the mode propagation and they have space reserved if desired/required/available.

The coupler used for the injection is a Thorlabs F220APC-1064, which allegedly collimates the beam from the fiber type we use to 2.4mm diameter, which I used as the target for the mode calculations. I coupled the first order diffracted beam to a ~60m fiber, which is a tad long but the only fiber I could locate that was long enough. The coupling efficiency from free-space to fiber is 47.5%, and we can currently get up to 63 mW out of the fiber.

Tomorrow Steve and I are going to pull the fiber through protective tubing and bring it to the AS port. The next step is then characterizing the beam out of the collimator to match it into the interferometer.

As far as the switching itself is concerned: I confirmed that the exponential decay is still present when looking at the fiber output. I located the DEI Pulser unit in the QIL lab, and also found several more AOMs, including a 200MHz Crystal Technologies one, same brand that the PSL has, where the ringdown was not observed. According to past elogs, with good polarizers we can expect an extinction ratio of ~200 from the Pockels cell, which should be fine, but it's going to be tradeoff switching speed <-> extinction (if the alternate AOM doesn't show this ringdown behavior).

Attachment 1: PSL_IR.pdf  531 kB  | Hide | Hide all
PSL_IR.pdf
Attachment 2: psl_aux_laser.pdf  3.740 MB  | Hide | Hide all
psl_aux_laser.pdf
ELOG V3.1.3-